OpenStudy (anonymous):

In isosceles triangle, vertex angle A is 5degrees more than one-half of base angle B. FIND THE SIZE OF EACH ANGLE OF THE TRIANGLE.

OpenStudy (anonymous):

vertex angle a = 40 degress and the base angles are 70 degrees each.

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

Let a be the vertex angle and b be either of the base angles. Solve the following simultaneous equations for a and b: {a = 5 + b/2, a + 2 b = 180} a = 40 and b = 70 40 + 2*70 = 180